# Determining unit groups and $\mathrm{K}_{1}$ of finite rings

### v1

@@ 8.2. Examples from noncommutative arithmetic orders
We consider computations of abelianizations of unit groups originating from [BHJ24] in the context of the stably free cancellation property of integral group rings and $\mathbf{Z}$-orders in general. The finite rings we consider arise as follow. For a $\mathbf{Z}$-order $\Lambda$ of a semisimple $\mathbf{Q}$-algebra $A$, any non-trivial decomposition $A\cong A_{1}\times A_{2}$ induces a fiber product involving a finite quotient ring $R$ of $\Lambda$. Of particular interest are integral group rings $\Lambda=\mathbf{Z}[G]$, where $G$ is related to binary polyhedral groups. In [BHJ24] the problem of determining $R^{\times}$ and $R^{\times\mathrm{ab}}$ was solved by constructing an explicit faithful permutation representation $\operatorname{Aut}_{\mathbf{Z}}(R^{+})\to\mathfrak{S}_{d}$, or (if $\operatorname{char}(R)=p$) a faithful linear representation $\operatorname{Aut}_{\mathbf{Z}}(R^{+})\to\operatorname{GL}_{n}(\mathbf{F}_{p})$. Together with the algorithms of [BB06, BHJ24] for determining generators of $R^{\times}$, one obtains a “realization” of $R^{\times}\to\operatorname{Aut}_{\mathbf{Z}}(R^{+})$ (considered as a black box group) as a subgroup of a direct product of permutation groups or linear groups over finite fields, from which $R^{\times\mathrm{ab}}$ can be determined. The algorithm of loc. cit. has been implemented using both OSCAR and Magma [BCP97]. Note that the complexity of this algorithm was not analyzed, but as the minimal degree of a faithful permutation presentation appears to grow fast, it is plausible that the algorithm is not polynomial in the length of $R$.
In the following, for $n\geq 2$ we denote by $\mathrm{Q}_{4n}$ the generalized quaternion group (also known as the dicyclic group) of order $4n$.

### v2

@@ 8.2. Examples from noncommutative arithmetic orders
We consider computations of abelianizations of unit groups originating from [BHJ24] in the context of the stably free cancellation property of integral group rings and $\mathbf{Z}$-orders in general. The finite rings we consider arise as follows. For a $\mathbf{Z}$-order $\Lambda$ of a semisimple $\mathbf{Q}$-algebra $A$, any non-trivial decomposition $A\cong A_{1}\times A_{2}$ induces a fiber product involving a finite quotient ring $R$ of $\Lambda$. Of particular interest are integral group rings $\Lambda=\mathbf{Z}[G]$, where $G$ is related to binary polyhedral groups. In [BHJ24] the problem of determining $R^{\times}$ and $R^{\times\mathrm{ab}}$ was solved by constructing an explicit faithful permutation representation $\operatorname{Aut}_{\mathbf{Z}}(R^{+})\to\mathfrak{S}_{d}$, or (if $\operatorname{char}(R)=p$) a faithful linear representation $\operatorname{Aut}_{\mathbf{Z}}(R^{+})\to\operatorname{GL}_{n}(\mathbf{F}_{p})$. Together with the algorithms of [BB06, BHJ24] for determining generators of $R^{\times}$, one obtains a “realization” of $R^{\times}\to\operatorname{Aut}_{\mathbf{Z}}(R^{+})$ (considered as a black box group) as a subgroup of a direct product of permutation groups or linear groups over finite fields, from which $R^{\times\mathrm{ab}}$ can be determined. The algorithm of loc. cit. has been implemented using both OSCAR and Magma [BCP97]. Note that the complexity of this algorithm was not analyzed, but as the minimal degree of a faithful permutation presentation appears to grow fast, it is plausible that the algorithm is not polynomial in the length of $R$.
In the following, for $n\geq 2$ we denote by $\mathrm{Q}_{4n}$ the generalized quaternion group (also known as the dicyclic group) of order $4n$. The examples provided in this section are also available on the author’s homepage.